# Mitigating Suboptimality of Deterministic Policy Gradients in Complex Q-functions

### v1

@@ G.2 Implementation Details
We used PyTorch (Paszke et al., 2019) for our implementation, and the experiments were primarily conducted on workstations with either NVIDIA GeForce RTX 2080 Ti, P40, or V32 GPUs on. Each experiment seed takes about 4-6 hours for Mine World, 12-72 hours for Mujoco, and 6-72 hours for RecSim, to converge. We use the Weights & Biases tool (Biewald, 2020) for plotting and logging experiments. All the environments were interfaced using OpenAI Gym wrappers (Brockman et al., 2016). We use the Adam optimizer (Kingma & Ba, 2014) throughout for training.
G.3 Common Hyperparameter Tuning
@@
To ensure fairness across all baselines and our methods, we performed hyperparameter tuning over parameters that are common across methods:
@@
  • Learning Rates of Actor and Critic: (Actor) We searched over learning rates $\{0.01,0.001,0.0001,0.0003\}$ and found that $0.0003$ was the most stable for the actor’s learning across all tasks. (Critic) Similar to the actor, we searched over the same set of learning rates and found the same value of $0.0003$ was the most stable for the critic’s learning across all tasks.
  • Network Sizes of Actor and Critic: For each task, we searched over simple 3 or 4 MLP layers to determine the network size that performed best but did not observe major differences. (Critic) To ensure a fair comparison, we used the same network size for the critic (Q-network) and surrogates across all methods within each task. (Actor) Similar to the critic, we used the same network size for the various actors in all the baselines and successive actors in SAVO within a particular task.
G.4 Hyperparameters
The environment and RL algorithm hyperparameters are described in Table 2.
Hyperparameter Mine World MuJoCo & Adroit RecSim
@@ G.4 Hyperparameters
Total Timesteps $10^{7}$ $3\times 10^{6}$ $10^{7}$
Number of epochs 5,000 8,000 10,000
@@
Buffer size $5\times 10^{5}$ $5\times 10^{5}$ $10^{6}$
Actor: LR $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
@@
Actor: $\epsilon$ decay steps $5\times 10^{6}$ $5\times 10^{5}$ $10^{7}$
@@
List Encoder LR $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
@@
Appendix H Q-Value Landscape Visualizations
H.1 1-Dimensional Action Space Environments
We analyzed the Q-value landscapes in Mujoco environments to show how successive critics help actors find better actions by reducing local optima. Figure 23 illustrates a typically smooth and easy-to-optimize Q-value landscape in unrestricted Inverted-Pendulum. Figure 24 illustrates that in restricted locomotion, the Q-value landscape (leftmost and rightmost figures) is uneven with many local optima. However, the Q-value landscapes learned by successive surrogates $\hat{\Psi}_{i}$ become successively smoother by removing local peaks below the Q-values of previously selected actions. This helps actors find closer to optimal actions than with a single critic.
Finally, when we plot the actions $a_{0},a_{1},a_{2}$ selected by the learned successive actors on the original Q-landscape (rightmost figure), we see they often achieve higher Q-values than $a_{0}$, the action a single actor has learned. Thus, the maximizer actor $\mu_{M}$ often finds closer to optimal actions than a single actor, resulting in better performance as shown in the return comparison between $\mu_{M}$ and single actor (Figure 11(c)) and the performance against baselines (Figure 8).
@@ Appendix H Q-Value Landscape Visualizations
Figure 23: Successive surrogate landscapes and the Q-landscape of Inverted Pendulum-v4.
Figure 24: Successive surrogate landscapes and Q landscape for Restricted Inverted-Pendulum and Restricted Inverted-Double-Pendulum environments.
H.2 High-Dimensional Action Space Environments
Figure 25 visualize Q-value landscapes for a TD3 agent in Hopper-v4. We project actions from the 3D action space of Hopper-v4 onto a 2D plane using Uniform Manifold Approximation and Projection (UMAP) and sample 10,000 actions evenly to ensure thorough coverage. The Q-values are plotted using trisurf, which may introduce some artificial roughness but offers more reliable visuals than grid-surface plotting. Despite limitations of dimensionality reduction — such as distortion of distances — the Q-landscape for Hopper-v4 reveals a large globally optimal region (shown in yellow), offering a clear gradient path that prevents the gradient-based actor from getting stuck in local optima.
In contrast, Hopper-Restricted (Figure 26) has more complex Q-landscapes due to valid actions being restricted in one of the hyperspheres shown in Figure 6. Consequently, these Q-landscapes appear to have more locally optimal regions than Hopper-v4. This creates many peaks where gradient-based actors might get trapped, degrading the resultant agent performance.
The curse of dimensionality limits conclusive analyses on higher dimensional environments like Walker2D-v4 (6D) and Ant-v4 (8D) because projecting to 2D causes significant information loss, making it difficult to assess convexity in their Q-landscapes.
Figure 25: Hopper-v4: Q landscape visualization at different states show a path to optimum.
Figure 26: Hopper-restricted: Q landscape visualization at different states show several local optima.

### v2

@@ 15.2 Implementation Details
We used PyTorch (Paszke et al., 2019) for our implementation, and the experiments were primarily conducted on workstations with either NVIDIA GeForce RTX 2080 Ti, P40, or V32 GPUs on. Each experiment seed takes about 4-6 hours for Mine World, 12-72 hours for Mujoco, and 6-72 hours for RecSim, to converge. We use the Weights & Biases tool (Biewald, 2020) for plotting and logging experiments. All the environments were developed using the OpenAI Gym Wrapper (Brockman et al., 2016). We use the Adam optimizer (Kingma & Ba, 2014) throughout.
15.3 Hyperparameters
The environment-specific and RL algorithm hyperparameters are described in Table 3.
15.4 Common Hyperparameter Tuning
@@
To ensure fairness across all baselines and our methods, We searched over hyper-parameters that are common across baselines;
@@
  • Learning rate of Actor and Critic: (Actor) We searched over $\{0.01,0.001,0.0001,0.0003\}$ and found that 0003 to be the most stable for the actor’s learning across all tasks. (Critic) Similarly to actor, we searched over $\{0.01,0.001,0.0001,0.0003\}$ and found that 0.0003 to be the most stable for the critic’s learning across all tasks.
  • Network Size of Actor and Critic: (Critic) In order for the fair comparison, we employed the same network size for the Q-network. We individually performed the architecture search on each task and found a specific network size performing the best in the task. (Actor) Similarly to critic, we employed the same network size for the actor components in the baseline and the cascading actors in SAVO. And, likewise, we performed the individual architecture search on each task and found a specific network size performing the best in the task.
Hyperparameter Mine World MuJoCo/Adroit RecSim
@@
Total Timesteps 10M 3M 10M
Number of epochs 5K 8K 10K
@@
Buffer size 500K 500K 1M
Actor: LR 0.0003 0.0003 0.0003
@@
Actor: $\epsilon$ decay steps 5M 500K 10M
@@
Critic: $\epsilon_{\text{start}}$ 1 1 1
Critic: $\epsilon_{\text{end}}$ 0.01 0.01 0.01
Critic: $\epsilon$ decay steps 500K 500K 2M
Critic: $\epsilon$ in Eval 0 0 0
@@
List Encoder LR 0.0003 0.0003 0.0003
@@
16 Scaling Number of Actors Needed in SAVO
Figure 26: SAVO optimized for number of actors against baselines. Comparison with baselines with SAVO optimized for the hyperparameter of the number of actors (10-15 actors) shows a more significant improvement than using only 3 actors in Figure 8.
16.1 Benchmarking SAVO with larger number of actors
While the results in main paper in Figure 8 use only 3 actors, we show in Figure 26 that SAVO’s improvement over TD3 and other baselines is even more significant when the number of actors is optimized and chosen as 10 (or 15 in RecSim).
17 Soft Actor-Critic (SAC): Mitigating Suboptimality with SAVO
We show that SAC (Haarnoja et al., 2018) is susceptible to gradient-descent-based local optima in the soft Q-landscape and demonstrate how SAVO improves performance when integrated with SAC.
SAC is susceptible to local optima in soft Q-landscape. DPG-based methods like TD3 optimize deterministic policies using:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{s\sim\rho^{\pi}}\left[Q^{\pi}(s,\pi(s))\right],$
where gradient ascent on $Q^{\pi}(s,\pi(s))$ often results in convergence to local optima due to the non-convexity of the Q-landscape.
SAC extends this framework by optimizing stochastic policies through entropy regularization, as:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{s\sim\rho^{\pi},a\sim\pi}\left[Q^{\pi}(s,a)+\alpha\mathcal{H}(\pi(\cdot|s))\right],$
where $\mathcal{H}(\pi(\cdot|s))=-\mathbb{E}_{a\sim\pi}[\log\pi(a|s)]$ is the entropy of the policy, weighted by $\alpha>0$.
However, despite the entropy-regularized objective, SAC’s actor is trained with gradient ascent on the soft Q-function $Q^{\pi}(s,a)$, which can be non-convex. Local optima in the (soft) Q-landscape arise from fundamental properties of the MDP and the non-convex relationship of actions and expected environment return. As a result, SAC policies are as prone to being trapped in local optima, in the KL-divergence sense, defined by the soft Q-landscape.
Figure 27: SAVO is complementary to TD3 and SAC. SAVO + SAC outperforms SAC in the three tasks evaluated: (i) Restricted Inverted Double Pendulum, (ii) Unrestricted Ant-v4, (iii) Unrestricted HalfCheetah-v4. SAVO improves or matches the performance of TD3 in the severely non-convex Q-landscape of the Restricted Inverted Double Pendulum and the high-dimensional action spaces of Ant-v4 and HalfCheetah-v4.
SAVO to mitigate SAC suboptimality. To address this challenge of SAC’s stochastic actor getting stuck in the soft Q-landscape’s local optima, we propose using SAVO as the actor architecture for SAC. In our approach, we introduce a maximizer stochastic actor $\pi_{M}$ that selects from successive stochastic actors $\nu_{i}(s;a_{<i})$ by maximizing:
$\pi_{M}(s):=\arg\max_{\nu_{0},...,\nu_{k}}\mathbb{E}_{s\sim\rho^{\pi},a\sim\pi}\left[Q^{\pi}(s,a)+\alpha\mathcal{H}(\pi(\cdot|s))\right].$
This SAC+SAVO approach leverages SAVO’s capacity to dynamically select policies that better navigate the soft Q-landscape while preserving SAC’s entropy-regularized exploration.
For this preliminary combination of SAC with SAVO, we do not employ the successive surrogates but only utilize successive actors with conditioning on previous actions.
Empirical Results. Figure 27 illustrates the relative performance of SAC, TD3, TD3+SAVO, and SAC+SAVO across the three tasks. Key findings include:
  • Hopper and Walker2D: SAC+SAVO significantly improves performance compared to SAC, demonstrating SAVO’s ability to overcome local optima in the soft Q-landscape.
  • Inverted Pendulum: SAC+SAVO exhibits faster convergence compared to SAC, further highlighting the synergy between SAVO and entropy-regularized stochastic policies.
  • Across all tasks, TD3+SAVO consistently outperforms TD3, confirming SAVO’s generalizability to deterministic policy optimization.
These results underscore the effectiveness of combining SAVO with both SAC and TD3, providing a robust solution to mitigate local optima and enhance exploration in complex control tasks.
18 Q-Value Landscape Visualizations
18.1 1-Dimensional Action Space Environments
We conducted a Q-space analysis across Mujoco environments to show that successive critics reduce local optima, aiding actors in optimizing actions. The outcomes are depicted in Figures 28 and 29.
Figure 28 illustrates a representative Q landscape from the easy environments, which are uniformly smooth. This uniformity in the primary Q space simplifies the identification of optimal actions.
Figure 29 shows that the primary Q landscape (leftmost and rightmost) in challenging environments is clearly uneven with several local optima. However, the Q landscapes learned by successive critics $Q_{i}$ demonstrate a gradual transition toward smoothness by pruning out the locally optimal peaks below the previously selected actions’ Q-values. This aids the actors in identifying improved actions that are better global optima over the primary critic. Finally, when visualized together on the primary critic (rightmost figure) the subsequent actions yield more enhanced Q-values than $a_{0}$, which would have been the action selected by a single actor.
@@ 18 Q-Value Landscape Visualizations
Figure 28: Successive Q landscape and primary Q landscape of Inverted Pendulum-v4.
Figure 29: Successive Q landscape and primary Q landscape across different Restricted Environments.
18.2 High-Dimensional Action Space Environments: Hopper-v4
In Fig.30 and Fig. 31, we visualize Q-landscapes for a TD3 agent across different environments, starting with Hopper-v4. Here, actions from the 3D action space are projected onto a 2D plane using UMAP, with 10,000 actions sampled at equal intervals to ensure adequate coverage. These Q-values are plotted using trisurf, introducing some artificial ruggedness but providing more reliable visualizations than grid-surface plotting. Despite the inherent limitations of dimensionality reduction—where the loss of one dimension distorts distances and relative positions—the Q-landscape for Hopper-v4 reveals a large globally optimal region (in yellow), offering a clear gradient path that minimizes the risk of the gradient-based actor getting stuck in local optima.
In Hopper-Restricted, the Q-landscapes become more complex due to the restriction of actions within a hypersphere, with suboptimal peaks where gradient-based actors can potentially get trapped. Although dimensional reduction limits conclusive analysis, these landscapes appear to have more local optima compared to Hopper-v4. For higher-dimensional environments like Walker2D-v4 (6D) and Ant-v4 (8D), projecting to 2D leads to significant information loss, making it difficult to assess convexity. Despite this, Walker2D-v4 shows a large optimal region where consecutive actions produce similar outcomes, indicating that contact-based tasks like Walker2D and Hopper do not inherently induce numerous local optima. However, for more complex environments like Ant-v4 and Walker2D-Restricted, the visualizations provide limited insights due to the challenges of dimensionality reduction.
Figure 30: Hopper-v4: Q landscape visualizations at different states show a path to optimum.
Figure 31: Hopper-restricted: Q landscape visualizations at different states show several local optima.